OpenStudy (anonymous):

How do you integrate x sin(2x) ?

OpenStudy (anonymous):

I would suggest integration by parts. \[\int\limits u dv=uv-\int\limits vdu\]

OpenStudy (anonymous):

yeah i chose u=xdx and du as dx . Then i picked dv=sin(2x) and i was trying to integrate that but idk what that is

OpenStudy (anonymous):

hmm, my integration is a little rusty, haven't had calculus for a year or so... let me see if I can work it.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

oh never mind i got it

OpenStudy (anonymous):

thanks though

OpenStudy (anonymous):

i realized i used the formula wrong

OpenStudy (anonymous):

haha no problem, glad you got it. Are you in calculus 1 or 2? 2 is pretty tough, getting good at integration is highly suggested.

OpenStudy (anonymous):

I'm takgin Advanced Placement Calc BC in high school

OpenStudy (anonymous):

yeah knowing how to integrate and differentiate is key in calculus...without those 2 skills its definitely hard

OpenStudy (anonymous):

Ahh, well good luck on that AP test haha, I remember it well

OpenStudy (anonymous):

oh man its coming up soooo soon and i feel soo unready for the test. it sucks becase my teacher was so awesome but i cant remeber things easily

OpenStudy (anonymous):

did u do well on the AP if u dont mind me asking